# Dark Energy from Time Crystals

### v1

@@ Abstract
In this work, we analyze a scalar field model which gives rise stable bound states in field space characterized by nonzero motion that breaks the underlying time translation symmetry of its Hamiltonian, known as time crystals. We demonstrate that an ideal fluid made up of these time crystals behaves as phantom dark energy characterized by an equation of state $w<-1$, speed of sound squared $c_{s}^{2}\geq 0$, and nonnegative energy density $\rho\geq 0$.
@@ I Introduction
A representative class of models with non-canonical kinetic terms give rise to the formation of crystalline type structures which oscillate in time rather than space, known as time crystals. They arise whenever a system described by a time- independent Hamiltonian has stable bound state solutions that spontaneously break the time translation symmetry of its Hamiltonian and reduce its time translation symmetry to a discreet subgroup that is periodic in time.
In this work, we investigate the time crystals model proposed in  [6, 7] and show that they behave as a ideal fluid with a phantom dark energy equation of state. Let’s start our investigation with a general non-canonical Hamiltonian in order to derive expressions for the energy density $\rho$, pressure $p$, equation-of-state parameter $w$, and speed of sound squared $c_{s}^{2}$ and then specialize our results to the time crystal model of  [6].
@@
where: $\phi$ is a scalar field and dot denotes the time derivative; $X=\dfrac{1}{2}\dot{\phi}^{2}>0$; $f(\phi)>0$ is a positive semi-definite function of $\phi$; $g(X)$ is a non trivial function of $X$, and $V(\phi)$ is the potential of the scalar field which we take to be a semi-definite positive function.
Below we first inspect the conditions required for an ideal fluid with a phantom equation of state $w<-1$, a nonnegative energy density $\rho\geq 0$, and a positive speed of sound squared $c_{s}^{2}\geq 0$. Then, we apply these considerations to the specific model of time crystals of  [6, 7] and explore its connection to phantom dark energy.
@@ II General Formalism
where $i$ runs over spatial indices.
The relevant Einstein equations that are sufficient for describing the ideal fluid are the Friedmann equation and the Bianchi identity. The Friedman equation is
@@
where $a(t)$ is the scale factor of the FRW metric, ˙$\dot{}=d/dt$. In general, we allow for other matter and energy components $\rho_{m}$ in the universe, not related to the non-canonical ideal fluid $\phi$. However, the phantom dark energy component soon dominates the expansion while the matter and radiation components dilute away, rendering them insignificant. For this reason we ignore the contribution of $\rho_{m}$ in the equations below.
@@
For an ideal fluid the Hamiltonian of the scalar field serves as the energy density of the cosmic fluid and the Lagrangian as the pressure term, Eqns. (4). Due to the non-canonical nature of the Lagrangian, the Lagrangian and the Hamiltonian expressions below are quite different from one another. Substituting this into Eq. (3) to obtain the Hamiltonian
$\rho$ $=\mathcal{H}=\Pi\dot{\phi}-\mathcal{L}=\dot{\phi}\left(f(\phi)g^{\prime}(X)\dot{\phi}\right)-\left(f(\phi)g(X)-V(\phi)\right)$
@@
$c_{s}^{2}=\dfrac{\partial p}{\partial\rho}=\dfrac{p_{,X}}{\rho_{,X}},$ (11)
where $p_{,X}=\dfrac{\partial p}{\partial X}$ and $\rho_{,X}=\dfrac{\partial\rho}{\partial X}$.
Computing $p_{,X}$ and $\rho_{,X}$ gives:
@@
Therefore, the expression for the speed of sound squared is
@@ II.1 Conditions for Stable Phantom Behavior with Positive Energy density
A phantom dark energy fluid has an equation-of-state parameter $w_{\phi}<-1$.
From Eq. (10), the condition $w_{\phi}<-1$ requires
@@
where $\rho+p$ using Eqs. (8) and (9)
$\rho+p$ $=2f(\phi)Xg^{\prime}(X).$ (16)
The condition for $w_{\phi}<-1$ then implies
@@
We further demand that the energy density of the ideal phantom dark energy fluid given by Eqn. (8) is non-negative. The requirement $\rho\geq 0$ yields
@@
Since $f(\phi)>0$, we can further simplify it as:
@@
This condition depends on the specific forms of $g(X)$, $V(\phi)$, and $f(\phi)$. However, given that $g^{\prime}(X)\leq 0$ for phantom dark energy ideal fluids, and further restrictions that we will discuss below in the context of time crystals, make the constraints on the potential energy term sufficiently weak to be accommodated by a large class of generic models. Potentials which satisfy the positive energy condition would only be weakly tuned.
If the speed of sound squared were negative we would be looking at a tachyonic fluid which can be unstable. Therefore, requiring the speed of sound squared of the fluid to be positive or zero provides the regions of stability in the field space.
The speed of sound squared is given by Eq. (14). The requirement of $c_{s}^{2}\geq 0$ leads to the following conditions on derivatives of $g(X)$
@@
Combining the above conditions ($c_{s}^{2}\geq 0$, $w<-1$, $\rho\geq 0$) for the phantom regime, we now compare these results against the time crystal model given in [6]. We show that the time crystal model of  [6, 7] gives rise to phantom dark energy and we seek to identify the ranges of $X$ where the phantom behavior is stable and the energy density is non-negative. This requires solving the inequalities described in this section simultaneously.
@@ III.1 A Time Crystals Lagrangian
Consider the Lagrangian density,
@@
proposed for time crystals in [6] where $\kappa_{2}>0$ and $\lambda_{2}>0$ are constants, $f(\phi)=1$, and $V(\phi)$ is a positive semi definite potential. In what follows, the non canonical kinetic term $g(X)$ of the previous section is therefore chosen to be that of the time crystal
@@
Based on these relations and Lagrangian equations, the meaning of $X_{t}=\kappa_{2}/6\lambda_{2}$ will soon become clear to define the turning points of the bound orbit of the time crystal that minimize energy, corresponding to $\dot{\phi}_{t}=\pm(\kappa_{2}/3\lambda_{2})^{1/2}\neq 0$. Importantly, there is motion $X\neq 0$ along the orbit $X_{t}$ which gives rise to the spontaneous breaking of the time translation symmetry of the Hamiltonian as is apparent from Eqn. (26). The effective potential $W(\phi)$ in the expression for the Hamiltonian, Eqn. (26), is given by
@@
Interestingly, note that there is an overall constant negative energy term $\Lambda_{T}$ in Eqn. (29) originating from time crystals contributing to the effective potential of the energy expression of Eqn. (26). The term $\Lambda_{T}$ which appears naturally from time crystals, can offset a large positive vacuum energy and alleviate the notorious fine tuning of problem of vacuum energy (by $123-$ orders of magnitude), through a far less tuned choice of its parameters $\lambda_{2},\kappa_{2}$. The parameters $\lambda_{2},\kappa_{2}$ which enter in a ratio, can take reasonable values and still be such that their combination in the time crystal contribution to vacuum energy $\Lambda_{T}$ is large or small enough to absorb the background’s positive vacuum energy.
@@
As we will see below, by writing the Hamiltonian and Lagrangian in this manner, both points $X_{t},X_{d}$ will play an important role in the dynamics of the fluid.
@@
$\frac{\partial\mathcal{L}}{\partial\phi}-3H\Pi=\frac{d}{dt}\frac{\partial\mathcal{L}}{\partial\dot{\phi}}=\frac{d}{dt}\Pi=\dot{\Pi}$ (33)
where $H=\frac{\dot{a}}{a}$ is the Hubble parameter.
These equations of motion are directly related to three points and regions of interest in the field space of the time crystals ideal fluid, that show up as $X_{t}$ and $X_{d}$ in Eqns.  (26) and  (30).
The requirement that kinetic energy is minimized leads to
$\dot{\Pi}=\ddot{\phi}(-\kappa_{2}+3\lambda_{2}\dot{\phi}^{2})=0$ (34)
which, for the nontrivial bound state $\dot{\phi}\neq 0$, implies that $\Pi(X)$ is a constant of motion along the turning points of time crystal orbit $\dot{\phi_{t}}=\pm\sqrt{\kappa_{2}/3\lambda_{2}}$ or $X=X_{t}$.
The turning points can be obtained by minimizing kinetic energy or equivalently from the cusps of the conjugate momentum. The cusps of the conjugate momentum $\Pi(X)$ indicate critical points in the fields space because the speed of sound diverges there $c_{s}^{2}->\infty$. these cusps are found by
@@
solutions of which occur exactly at the turning points of the time crystal $\dot{\phi}_{t}=\pm\sqrt{\kappa_{2}/3\lambda{2}}$ or $X=X_{t}$. The latter, in combination with the fact that the speed of sound squared diverges at the boundary, indicate that these bound structures, the time crystals, are stable against perturbations and the field is confined to the time crystal region $X\leq X_{t}$. Hence, the time crystal bound state is a state of minimum energy, despite that the underlying time translation symmetry is broken and there is motion $X\neq 0$ along the orbit. The meaning of ($X_{t},\pm\phi_{t}$) then becomes clear as the turning points that define the orbit of the bound state, but which, due to motion break time translation symmetry of the Hamiltonian and become time crystals.
Another point of interest is the one where the group velocity of the field vanishes and the kinetic energy is once again minimized, the point where $\Pi=\frac{\mathcal{L}}{\partial\dot{\phi}}=0$. As can be seen from Eq. (28), this occurs at $X=X_{d}=\frac{\kappa_{2}}{2\lambda_{2}}$ or equivalently along the orbit defined by the special points $\phi_{d}=\pm\sqrt{\kappa_{2}/\lambda_{2}}$ that enter the Lagrangian expression, Eqn. (30).
@@ III.2 Phantom Dark Energy from the Time Crystal model
The condition for Phantom dark energy behavior, ($w<-1$), from Eq. (18) is: $g^{\prime}(X)<0$. Substituting $g^{\prime}(X)$ from Eq. (36):
$-\kappa_{2}+2\lambda_{2}X<0.$ (38)
and solving for $X$:
@@
Hence, phantom behavior occurs anywhere in the region below the point $X=X_{d}$ where group velocity of the field and conjugate momentum $\Pi$ vanish:
@@
The condition for nonnegative energy density ($\rho\geq 0$), with $f(\phi)=1$ is given by:
@@
For potentials that satisfy $W(\phi)\geq 0$, equivalently $V(\phi)\geq\frac{\kappa_{2}^{2}}{12\lambda_{2}}$, the energy density is positive everywhere, as per Eqn.  (26).
The general case of $V(\phi)$ in the above equation is a quadratic inequality in $X$ which can be solved for a range of $X$ where $\rho\geq 0$.
Given the bounds on $g(X),W(\phi)$ it is straightforward to check that the energy density is always positive in the region of interest where the ideal fluid as phantom behavior, which included the time crystal $X\leq X_{t}$.
The equation of state parameter $w_{\phi}$ for the ideal fluid described by the time crystals model is given by
@@
for $X\leq X_{d}$.
Finally let’s check the condition for stability ($c_{s}^{2}\geq 0$) of the phantom fluid.
From Eq. (14), we have:
@@
Define:
$N$ $=-\kappa_{2}+2\lambda_{2}X,$ (45)
$D$ $=-\kappa_{2}+6\lambda_{2}X.$ (46)
Let’s consider the regions:

### v2

@@ Abstract
In this work, we analyze a scalar field model which gives rise to stable bound states in field space characterized by nonzero motion that breaks the underlying time translation symmetry of its Hamiltonian, known as time crystals. We demonstrate that an ideal fluid made up of these time crystals behaves as phantom dark energy characterized by an equation of state $w<-1$, speed of sound squared $c_{s}^{2}\geq 0$, and nonnegative energy density $\rho\geq 0$.
@@ I Introduction
A representative class of models with non-canonical kinetic terms give rise to the formation of crystalline type structures which oscillate in time rather than space, known as time crystals. They arise whenever a system described by a time- independent Hamiltonian has stable bound state solutions that spontaneously break the time translation symmetry of its Hamiltonian and reduce its time translation symmetry to a discreet subgroup that is periodic in time. In this work, we investigate the time crystals model proposed in  [6, 7] and show that it behaves as a ideal fluid with a phantom dark energy equation of state.
Starting our investigation with a general non-canonical Lagrangian, we first derive the general expressions for the energy density $\rho$, pressure $p$, equation-of-state parameter $w$, and speed of sound squared $c_{s}^{2}$, and then apply these results to the time crystal model of  [6].
@@
where: $\phi=\phi_{0}(t)+\delta\phi(t,x)$ is a scalar field with $\phi_{0}$ the background field and $\delta\phi$ perturbations about it. As is typical of inflationary models, the background field $\phi_{0}(t)$ is homogeneous while perturbations are functions of space and time ($x,t$). The kinetic term defined as $X=\dfrac{1}{2}g_{\mu\nu}\nabla{\phi}^{\mu}\nabla^{\nu}$ reduces to $X=\frac{1}{2}\dot{\phi}^{2}>0$ for the homogeneous mode $\phi_{0}$. Here, $f(\phi)>0$ is a positive semi-definite function of $\phi$; $g(X)$ is a non trivial function of $X$, and $V(\phi)$ is the potential of the scalar field which we take to be a semi-definite positive function.
The conditions required for an ideal fluid with a phantom equation of state $w<-1$, a nonnegative energy density $\rho\geq 0$, and a positive speed of sound squared $c_{s}^{2}\geq 0$ are reviewed in Section II. In Section III, we apply these considerations to the specific model of time crystals of  [6, 7] and explore its connection to phantom dark energy. From here on, we focus on the dynamics of the homogeneous field $\phi_{0}$ and drop the subscript 0. The study of perturbations is reported in a companion paper  [10].
@@ II General Formalism
where $i$ runs over spatial indices. Hence, for an ideal fluid its pressure $p$ and energy density $\rho$ are identified with the Lagrangian and Hamiltonian of the scalar field respectively.
The relevant Einstein equations are the Friedmann equation and the Bianchi identity. The Friedman equation is
@@
where $a(t)$ is the scale factor of the FRW metric with line element $ds^{2}=-dt^{2}+a(t)^{2}dx_{(3)}^{2}$, and ˙$\dot{}=d/dt$. In general, we could allow for other matter and energy components $\rho_{m}$ in the universe, not related to the non-canonical ideal fluid $\phi$. However, the phantom dark energy component soon dominates the expansion while the matter and radiation components dilute away, rendering them insignificant. For this reason we ignore the contribution of $\rho_{m}$ in the equations below.
@@
Due to the non-canonical nature of the Lagrangian, the Lagrangian and the Hamiltonian expressions below are quite different from one another. Substituting this into Eq. (3) to obtain the Hamiltonian
$\rho=\mathcal{H}=\Pi\dot{\phi}-\mathcal{L}$ (8)
@@
$c_{s}^{2}=\dfrac{p_{,X}}{\rho_{,X}},$ (12)
where $p_{,X}=\dfrac{\partial p}{\partial X}$ and $\rho_{,X}=\dfrac{\partial\rho}{\partial X}$, and,
@@
Therefore, the expression for the speed of sound squared becomes
@@ II.1 Conditions for Stable Phantom Behavior with Positive Energy density
A phantom dark energy fluid has an equation-of-state parameter $w_{\phi}<-1$. We demand that the energy density of the phantom fluid given by Eqn. (9) is positive, and explore the field range under which the three quantities of interest simultaneously satisfy their constraints ($\rho\geq 0,c_{s}^{2}\geq 0,w_{\phi}\leq-1$) .
Given the requirement that $\rho\geq 0$, from Eqn. (11) the condition $w_{\phi}<-1$ is equivalent to
@@
Using Eqns. (9) and (10), the condition $w_{\phi}<-1$ then implies
@@
The requirement $\rho\geq 0$ is equivalent to
@@
Since $f(\phi)>0$, it can be further simplified as:
@@
an inequality which can be solved, given a specific expression of $g(X)$ and $V(\phi)$. The $\rho\geq 0$ requirement of Eqn. (20) depends on the specific forms of $g(X)$, $V(\phi)$, and $f(\phi)$.
However, given that $g^{\prime}(X)\leq 0$ for phantom dark energy ideal fluids, and further restrictions that we will discuss below in the context of time crystals, that limit the field from exploring a wide range of the field space and confine it to the crystal region, as well as the coupling of the field to gravity on an expanding FRW background, make the dependence on the potential energy term sufficiently weak to be accommodated by a large class of generic models. Potentials which satisfy the positive energy condition would only be weakly tuned.
If the speed of sound squared were negative we would be looking at a tachyonic fluid which is wildly unstable. Therefore, requiring the speed of sound squared of the fluid to be positive or zero provides the regions of stability in the field space.
The speed of sound squared is given by Eq. (15). Our third requirement that $c_{s}^{2}\geq 0$ leads to the following constraint on derivatives of $g(X)$
@@
Combining the three constraints ($c_{s}^{2}\geq 0$, $w<-1$, $\rho\geq 0$) for the phantom regime, we now apply them to the time crystal model given in [6]. We show that the time crystal model of  [6, 7] gives rise to phantom dark energy and the range of $X\leq X_{t}$ inside the crystal boundaries satisfy all three constraints. To demonstrate this, it requires solving the inequalities described in this section simultaneously. Notice that, in contrast to k-essence models where the field can explore a large range in field space, the case of time crystals the field becomes more intriguing since the field is confined to a small range within the boundaries of the crystal, and forbidden to climb out of the crystal range.
@@ III.1 A Time Crystals Lagrangian
Consider the time crystal Lagrangian density of  [6],
@@
where $\kappa_{2}>0$ and $\lambda_{2}>0$ are constants, $f(\phi)=1$, and $V(\phi)$ is a positive semi definite potential. The non canonical kinetic term $g(X)$ of the previous section is
@@
$=\dot{\phi}(-\kappa_{2}+\lambda_{2}\dot{\phi}^{2})$ (29)
The boundary $X_{t}=\kappa_{2}/6\lambda_{2}$ defines the turning points of the bound orbit of the time crystal that minimize kinetic energy $\frac{\partial\mathcal{H}}{\partial X}|_{X=X_{t}}=0$, corresponding to $\dot{\phi}_{t}=\pm(\kappa_{2}/3\lambda_{2})^{1/2}\neq 0$. Note that there is motion $X\neq 0$ along the orbit $X_{t}$ which gives rise to the spontaneous breaking of the time translation symmetry of the Hamiltonian as is apparent from Eqn. (26). The effective potential $W(\phi)$ in the expression for the Hamiltonian, Eqn. (26), is given by
@@
Rewriting the energy density in this manner, illustrates the fact that the crystal contributes an overall constant negative energy term $\Lambda_{T}$ in Eqn. (30) that contributes to the effective potential of the energy expression of Eqn. (26). The term $\Lambda_{T}$ which appears naturally from time crystals, can offset a large positive vacuum energy and alleviate the notorious fine tuning of problem of vacuum energy (by $123-$ orders of magnitude), through a far less tuned choice of its parameters $\lambda_{2},\kappa_{2}$. The parameters $\lambda_{2},\kappa_{2}$ which enter in a ratio, can take reasonable values and still be such that their combination in the time crystal contribution to vacuum energy $\Lambda_{T}$ is large or small enough to absorb the background’s positive vacuum energy.
@@
where $X_{d}$ is defined as the locus of the Lagrangian $\frac{\partial\mathcal{L}}{\partial X}|_{X=X_{d}}=0$. As we will see below, by writing the Hamiltonian and Lagrangian in this manner, both points $X_{t},X_{d}$ will play an important role in the dynamics of the fluid.
@@
$\frac{1}{a(t)^{3}}\frac{da(t)^{3}\Pi(X)}{dt}=\frac{\partial\mathcal{L}}{\partial\phi}$ (34)
which, when expanded reads $\frac{\partial\mathcal{L}}{\partial\phi}=3H\Pi+\dot{\Pi}$, with $H=\frac{\dot{a}}{a}$ is the Hubble parameter. The equation of motion can be written specifically in terms of the field, by factoring out $\rho_{X}$ and using the definition of the speed of sound, as
$\ddot{\phi}+3Hc_{s}^{2}\dot{\phi}+\frac{1}{\rho_{X}}\frac{\partial V(\phi)}{\partial\phi}=0$ (35)
The equations of motion are directly related to three points and regions of interest in the field space of the time crystals ideal fluid, that show up as $X_{t}$ and $X_{d}$ in Eqns.  (26) and  (31).
The set of Einstein’s equations, e.g. $H\simeq\sqrt{\rho}$, coupled to the field equation of motion Eqn. (35), and the perturbations equation  [10] can be solved numerically and we will report these results in an upcoming paper. However, since the Hubble parameter grows and the scale factor diverges within a finite time interval, according to Eqn. (34) the conjugate momentum is driven to zero as $\Pi\simeq\frac{1}{a^{3}}$ during the same time interval $t\simeq t_{0}w/(1+w)$ with $t_{0}$ the present time  [9], as $a(t)\rightarrow\infty$.
The turning points can be obtained by minimizing kinetic energy or equivalently from the cusps of the conjugate momentum. The cusps of the conjugate momentum $\Pi(X)$ indicate critical points in the fields space because, as can be seen by a direct inspection, the speed of sound diverges there $c_{s}^{2}->\infty$. These cusps are found by
@@
solutions of which occur exactly at the turning points of the time crystal $\dot{\phi}_{t}=\pm\sqrt{\kappa_{2}/3\lambda{2}}$ or $X=X_{t}$. The latter, in combination with the fact that the speed of sound squared diverges at the boundary, indicate that these bound structures, the time crystals, are stable against perturbations  [10] and the field is confined to the time crystal region $X\leq X_{t}$ where the boundary of the crystal behaves as a hard wall. The latter should be contrasted to k-essence models where the field explores a large range in field space. Hence, the time crystal bound state is a state of minimum kinetic energy, despite that the underlying time translation symmetry is broken and there is motion $X\neq 0$ along the orbit. The meaning of ($X_{t},\pm\phi_{t}$) then becomes clear as the turning points that define the orbit of the bound state, but which, due to motion break time translation symmetry of the Hamiltonian and become time crystals.
Another point of interest is the one where the group velocity of the field vanishes and the kinetic energy is once again minimized, the locus of the Lagrangian $\Pi=\frac{\mathcal{L}}{\partial\dot{\phi}}=0$. As can be seen from Eq. (29), this occurs at $X=X_{d}=\frac{\kappa_{2}}{2\lambda_{2}}$ or equivalently along the orbit defined by the special points $\phi_{d}=\pm\sqrt{\kappa_{2}/\lambda_{2}}$ that enter the Lagrangian expression, Eqn. (31). However these points lie outside the crystal’s allowed field range, $X_{d}>X_{t}$. Due to the fact that the boundary of the crystal $X=X_{t}$ is a hard wall type with a diverging speed of sound squared, the field cannot access regions $X>X_{t}$ beyond the crystal boundary, therefore it can not reach $X_{d}$.
@@ III.2 Phantom Dark Energy from the Time Crystal model
The condition for Phantom dark energy behavior, ($w<-1$) provided that $\rho\geq 0$, Eq. (18), is: $g^{\prime}(X)<0$. Substituting $g^{\prime}(X)$ from Eq. (37) gives
@@
Hence, phantom behavior occurs anywhere in the region below the point $X\leq X_{d}$ where group velocity of the field and conjugate momentum $\Pi$ vanish:
@@
The requirement for a nonnegative energy density ($\rho\geq 0$), with $f(\phi)=1$ is satisfied by:
@@
which, for any specified potential, is straightforward to solve as a quadratic inequality. For potentials that satisfy $W(\phi)\geq 0$, equivalently $V(\phi)\geq\frac{\kappa_{2}^{2}}{12\lambda_{2}}$, the energy density is positive everywhere, as it’s clear from Eqn.  (26).
Given the above bounds on $g(X),W(\phi)$ it is easy to see that the energy density is always positive in the region of interest where the ideal fluid as phantom behavior, which included the time crystal $X\leq X_{t}$.
The equation of state parameter $w_{\phi}$ for the ideal fluid described by the time crystals model is of a phantom type,
@@
for all $X\leq X_{d}$.
Finally let’s check the condition for stability ($c_{s}^{2}\geq 0$) of the phantom fluid that enters the field equation, Eqn. (35). At this point, the behavior of the model becomes interesting.
Using Eq. (15), we have:
@@
which varies across the $X-$space.
The variation of the behavior of the field and of the speed of sound squared across the whole range of ($\phi,X$) divides this range into three regions, as summarized below. Given:
$p_{X}$ $=-\kappa_{2}+2\lambda_{2}X,$ (45)
$\rho_{X}$ $=-\kappa_{2}+6\lambda_{2}X.$ (46)
The three regions are: